# The Law of Multi-Model Collaboration: Scaling Limits of Model Ensembling for Large Language Models

### v1

@@ 5.1 Multi-model Collaboration as an Extension of Scaling Laws
Classical scaling laws characterize how the performance of a single model improves as a function of its parameter count, data, and compute. Implicit in these laws is the assumption that scaling occurs along a single architectural and representational trajectory (Kaplan et al., 2020; Hoffmann et al., 2022). Our results indicate that multi-model collaboration is a qualitatively distinct kind of scaling.There is a stable power law relationship between oracle loss and aggregated parameter budget for multi-model system which means that multi-model system has its own scaling regime. This regime is not just a simple pile-up of single-model behavior. On the contrary, better scaling exponents and lower asymptotic loss floors show that working together lets us reach parts of the performance frontier that we can’t get to just by making one model bigger. And so this leads us to think of the total parameter budget of a pool of models as a new kind of scaling variable that’s somewhat similar to the number of parameters for a single model, but it’s living in a much bigger model space.
@@ 5.2 Interpretation of Scaling Exponents
The scaling exponent $\alpha$ reflects the efficiency with which additional parameters translate into performance gains. The consistently larger exponents observed for multi-model systems indicate that parameters distributed across multiple models can be utilized more efficiently than parameters concentrated within a single model. Notably, homogeneous multi-model ensembles show extremely steep scaling exponents for small parameter budgets, indicating quick early improvements.But it has early saturation too, with asymptotic loss floors that match those of a single model.This pattern implies that merging similar models mainly reduces redundancy or variance within a common inductive bias, rather than substantially enlarging representational capacity.In contrast, heterogeneous ensembles get better scaling exponents and much smaller asymptotic loss floors.This combination implies that architectural and training diversity enables sustained gains across scales, consistent with the notion that different models encode complementary hypotheses about the data distribution (Dietterich, 2000).
@@ 5.3 Lower Loss Floors and the Limits of Single-Model Scaling
The reduction of the asymptotic loss floor that we see with heterogeneous multi-model systems is significant for our understanding of the limits on scaling single models.In classical scaling laws, the loss floor is often interpreted as reflecting irreducible error under a fixed model class and training regime (Kaplan et al., 2020; Henighan et al., 2020).From our findings we can conclude that it’s not just the number of parameters which determines this floor, but also the variety of inductive biases at the disposal of the system. By allowing multiple model families to contribute, multi-model collaboration effectively expands the hypothesis space, thereby lowering the attainable error bound (Domingos, 2012). So from this point of view, the loss floor for single-model scaling laws needs to be understood as being tied to some particular lineage of architecture, rather than as some sort of hard stop.
@@ 6 Limitations and Scope of the Law
While the empirical results presented in this work provide strong evidence for the Law of Multi-model Collaboration, it is important to clarify the scope of the law and acknowledge its limitations.These limitations arise from the theoretical perspective taken here, and they do not detract from the main contribution, which is identifying the basic performance limits rather than to suggest practical systems.
@@ 6.1 Idealized Oracle Assumption
An important abstraction in our formulation is the oracle ensemble, which selects, for each input text, the model within a pool that achieves the lowest loss.It’s an intentionally idealized assumption that does not correspond to any feasible inference time mechanism.In practice, it is usually impossible to determine which model performs best for each input without knowing the ground-truth loss.However, this abstraction serves a role analogous to idealized assumptions commonly employed in prior scaling-law analyses, such as optimal data usage or perfect optimization.The oracle ensemble sets an upper limit on the performance that can be achieved with multi-model collaboration, thus enabling the study of intrinsic scaling behavior without considering algorithmic inefficiency. Therefore our results should be seen as theoretical limits rather than deployable architectures.
@@ 6.2 Upper Bounds versus Realizable Systems
The Law of Multi-model Collaboration describes the behavior of Pareto-optimal lower envelopes under an oracle assumption. As such, it does not claim that practical multi-model systems can attain the same performance. Real world implementations have other constraints such as routing errors, coordination overhead, latency, and compute costs.Nonetheless, upper-bound analyses play a critical role in guiding system design by revealing what is possible in principle. Like how single-model scaling laws give us ideas on what kind of models to use and where to put the data even though they make some idealized assumptions, this law gives us something to compare against and encourage people to work on better ways of putting different models together.
6.3 Restriction to Pairwise Model Combinations
Our empirical analysis focuses on pairwise combinations of models. This restriction is due to computational constraints; the number of potential subsets increases exponentially as the size of the model pool grows.Higher order combinations may give even better results, but the scaling behavior is already evident for pairs. Importantly, the law itself is not inherently limited to pairs. The formulation in Section 2 applies to arbitrary model subsets, and extending the empirical analysis to larger ensembles is a natural direction for future work.Higher-order combinations are expected to make the observed trends more pronounced, possibly resulting in even lower performance frontiers.
@@ 6.4 Dependence on the Model Pool
The fitted scaling parameters rely on the makeup of the underlying model pool.Different choices for model families, parameter ranges, or training protocols can cause differences in scaling exponents and loss floors.Accordingly, the numerical values reported in this work should not be interpreted as universal constants.Rather, the law asserts the existence of a stable power-law relationship and systematic differences between single-model and multi-model scaling regimes. The qualitative conclusions about enhanced scaling efficiency and decreased asymptotic loss with heterogenous collaboration hold up well for all kinds of models.
@@ 6.5 Loss Definition and Tokenization Effects
To make it possible for different models with different tokenizers to have a fair comparison, we evaluate the text-level loss by adding up all the token-level cross entropy values instead of taking their average. This choice avoids the normalization artifacts from tokenizer-dependent sequence lengths, but it means that the loss values depend on how we tokenize things.While this formulation is well-suited for comparing oracle performance bounds, alternative loss normalizations may be more appropriate in other contexts. The law itself is independent from the particular choice of the loss function, and it will be interesting for future work to study the effect of different evaluation metrics on the observed scaling behavior.
@@ 7 Conclusion
We introduce the Law of Multi-model Collaboration, a scaling law defining the theoretical performance limits of systems composed of multiple pretrained LLMs. By employing an oracle-based, method-agnostic formulation, we isolate the intrinsic behavior of multi-model collaboration and provide a principled comparison to classical single-model scaling laws. Our empirical study across diverse open-source models demonstrates that multi-model systems follow stable power-law scaling relative to aggregated parameter budgets. We have conducted a systematic empirical study on various open-source language models, which shows that multi-model systems show stable power-law scaling regarding their combined parameter budget. Compared to single-model scaling, multi-model collaboration can achieve better scaling efficiency, and in heterogenous pools, it has much lower asymptotic loss floors.These results suggest that distributing parameters across multiple models—especially those with diverse inductive biases—can be more effective than monolithic scaling within a single architecture. Beyond the empirical results, this work reframes multi-model systems as objects governed by regular and predictable scaling behavior rather than as ad-hoc engineering solutions.The proposed law identifies model diversity as a fundamental scaling dimension, alongside parameters, data, and compute. From this perspective, the perceived limits of single-model scaling appear to be conditional on architectural lineage rather than absolute boundaries. While the Law of Multi-model Collaboration defines theoretical upper bounds rather than realizable systems, such bounds are essential for clarifying what is achievable in principle. We hope this work inspires further research into diversity-aware scaling, principled integration mechanisms, and broader theories of collective intelligence in machine learning.

### v2

@@ 5.1 Multi-model Collaboration as an Extension of Scaling Laws
Classical scaling laws characterize how the performance of a single model improves as a function of its parameter count, data, and compute. Implicit in these laws is the assumption that scaling occurs along a single architectural and representational trajectory (Kaplan et al., 2020; Hoffmann et al., 2022). Our results indicate that multi-model collaboration is a qualitatively distinct kind of scaling. There is a stable power law relationship between oracle loss and aggregated parameter budget for a multi-model system, which means that the multi-model system has its own scaling regime. This regime is not just a simple pile-up of single-model behavior. On the contrary, better scaling exponents and lower asymptotic loss floors show that working together lets us reach parts of the performance frontier that we can’t get to just by making one model bigger. And so this leads us to think of the total parameter budget of a pool of models as a new kind of scaling variable that’s somewhat similar to the number of parameters for a single model, but it’s living in a much bigger model space.
@@ 5.2 Interpretation of Scaling Exponents
The scaling exponent $\alpha$ reflects the efficiency with which additional parameters translate into performance gains. Notably, the scaling exponents for single-model and multi-model ensemble frontiers are remarkably similar ($\alpha=0.3578$ and $\alpha=0.3502$, respectively), indicating that the rate of loss reduction with respect to parameter count follows a consistent power-law across both regimes. This similarity suggests that the fundamental efficiency of parameter utilization, in terms of marginal gains per additional parameter, is comparable whether parameters are concentrated within a single architecture or distributed across multiple models. However, the critical distinction lies not in the scaling rate, but in the baseline performance level from which scaling occurs. Multi-model collaboration effectively shifts the entire performance curve downward, achieving lower loss at every parameter budget. This vertical shift manifests most dramatically in the asymptotic regime, where the multi-model ensemble approaches a substantially lower loss floor than any single model. In the pairwise diversity analysis, we observe a revealing trade-off between scaling rate and asymptotic performance. Homogeneous same-family pairs exhibit a steep scaling exponent ($\alpha=0.8501$), reflecting rapid early gains at small parameter budgets, but quickly saturate at an asymptotic loss floor ($L_{\infty}=2.2314$) comparable to single models ($L_{\infty}=2.21$). This suggests that combining architecturally similar models primarily reduces variance or redundancy within a shared inductive bias without fundamentally expanding representational capacity. In contrast, heterogeneous cross-family pairs exhibit a more moderate scaling exponent ($\alpha=0.5516$) but achieve a substantially lower asymptotic loss floor ($L_{\infty}=2.0377$). This pattern is consistent with the notion that different model families encode complementary hypotheses about the data distribution (Dietterich, 2000). The steeper exponent of homogeneous pairs indicates efficient exploitation of redundancy, while the lower floor of heterogeneous pairs reveals the value of architectural diversity in accessing distinct regions of the hypothesis space.
@@ 5.3 Lower Loss Floors and the Limits of Single-Model Scaling
The 43% reduction in asymptotic loss floor observed for multi-model ensembles (from 2.21 to 1.25) has profound implications for our understanding of the limits of single-model scaling. In classical scaling laws, the loss floor is often interpreted as reflecting irreducible error under a fixed model class and training regime (Kaplan et al., 2020; Henighan et al., 2020). Our findings demonstrate that this floor is not an absolute boundary but rather a constraint specific to individual architectural lineages. The dramatic reduction in loss floor reveals that multi-model collaboration fundamentally expands the hypothesis space accessible to the system. While a single model, regardless of its size, is constrained by the inductive biases inherent in its architecture and training procedure, an oracle ensemble can adaptively leverage complementary strengths across multiple architectures. This suggests that the variety of inductive biases available to the system not merely the total number of parameters determines the achievable performance limit. From this perspective, the loss floor for single-model scaling laws should be understood as architecture-specific rather than representing a fundamental limit of the modeling task itself. Multi-model collaboration effectively transcends this constraint by accessing a broader region of the hypothesis space (Domingos, 2012), enabling performance levels that are structurally inaccessible to any individual architecture.
@@ 6 Limitations and Scope of the Law
While the empirical results presented in this work provide strong evidence for the Law of Multi-model Collaboration, it is important to clarify the scope of the law and acknowledge its limitations. These limitations arise from the theoretical perspective taken here, and they do not detract from the main contribution, which is identifying the basic performance limits rather than suggesting practical systems.
@@ 6.1 Idealized Oracle Assumption
An important abstraction in our formulation is the oracle ensemble, which selects, for each input text, the model within a pool that achieves the lowest loss. It is an intentionally idealized assumption that does not correspond to any feasible inference time mechanism. In practice, it is usually impossible to determine which model performs best for each input without knowing the ground-truth loss. However, this abstraction serves a role analogous to idealized assumptions commonly employed in prior scaling-law analyses, such as optimal data usage or perfect optimization. The oracle ensemble sets an upper limit on the performance that can be achieved with multi-model collaboration, thus enabling the study of intrinsic scaling behavior without considering algorithmic inefficiency. Therefore, our results should be seen as theoretical limits rather than deployable architectures.
@@ 6.2 Upper Bounds versus Realizable Systems
The Law of Multi-model Collaboration describes the behavior of Pareto-optimal lower envelopes under an oracle assumption. As such, it does not claim that practical multi-model systems can attain the same performance. Real-world implementations have other constraints, such as routing errors, coordination overhead, latency, and compute costs. Nonetheless, upper-bound analyses play a critical role in guiding system design by revealing what is possible in principle. Like how single-model scaling laws give us ideas on what kind of models to use and where to put the data, even though they make some idealized assumptions, this law gives us something to compare against and encourages people to work on better ways of putting different models together.
6.3 Pareto-Pruning as an Approximation Strategy
To address the combinatorial explosion inherent in exploring all $2^{71}-1$ possible model subsets, we employ an iterative Pareto-pruning strategy that constructs size-$k$ ensembles by incrementally adding models to Pareto-optimal ensembles of size k1k-1. While this greedy heuristic enables tractable exploration of the multi-model scaling frontier, it does not guarantee exhaustive coverage of all potentially optimal configurations. Specifically, it is possible that some non-Pareto-optimal ensembles of size k1k-1 could, when augmented with appropriate models, yield Pareto-optimal ensembles of size $k$ that are missed by our procedure. Nonetheless, the resulting 3,146 Pareto-optimal configurations exhibit stable and predictable power-law scaling behavior, suggesting that the pruning strategy captures the fundamental scaling trends. The close agreement between the empirical frontier and the fitted scaling law provides evidence that the explored configurations are representative of the broader multi-model scaling regime. Future work could investigate alternative search strategies, such as genetic algorithms or reinforcement learning-based exploration, to more comprehensively sample the multi-model configuration space.
@@ 6.4 Dependence on the Model Pool
The fitted scaling parameters rely on the makeup of the underlying model pool. Different choices for model families, parameter ranges, or training protocols can cause differences in scaling exponents and loss floors. Accordingly, the numerical values reported in this work should not be interpreted as universal constants. Rather, the law asserts the existence of a stable power-law relationship and systematic differences between single-model and multi-model scaling regimes. The qualitative conclusions about enhanced scaling efficiency and decreased asymptotic loss with heterogeneous collaboration hold up well for all kinds of models.
@@ 6.5 Loss Definition and Tokenization Effects
To enable fair comparison across models with heterogeneous tokenizers, we compute the text-level loss as the sum of token-level cross-entropy values, which is then normalized by the average token sequence length across all models and all test texts. This formulation preserves sensitivity to modeling efficiency at the text level while accounting for tokenization granularity differences. The normalization by a shared average sequence length ensures that models are evaluated on a consistent scale, avoiding biases introduced by tokenizer-specific sequence lengths. While this formulation is well-suited for comparing oracle performance bounds, alternative loss normalizations may be more appropriate in other contexts. The law itself is independent of the particular choice of the loss function, and it will be interesting for future work to study the effect of different evaluation metrics on the observed scaling behavior.
@@ 7 Conclusion
We introduce the Law of Multi-model Collaboration, a scaling law defining the theoretical performance limits of systems composed of multiple pretrained LLMs. By employing an oracle-based, method-agnostic formulation, we isolate the intrinsic behavior of multi-model collaboration and provide a principled comparison to classical single-model scaling laws.
Through systematic exploration of model combinations ranging from 2 to 71 models via iterative Pareto-pruning, we demonstrate that multi-model systems follow stable power-law scaling with respect to aggregated parameter budgets. While exhibiting similar scaling exponents to single models ($\alpha\approx 0.35$), multi-model ensembles achieve a dramatic 43% reduction in asymptotic loss floor (from 2.21 to 1.25), fundamentally shifting the achievable performance frontier downward. This indicates that multi-model collaboration does not merely accelerate convergence along the single-model trajectory but rather enables access to performance regimes that are structurally inaccessible to any individual architecture.
Additionally, pairwise analysis reveals that model diversity is a critical driver of collaboration gains: heterogeneous cross-family ensembles consistently outperform homogeneous same-family ensembles under equivalent parameter budgets, achieving substantially lower asymptotic loss floors. These results suggest that distributing parameters across multiple models—especially those with diverse architectural and training characteristics—can be more effective than monolithic scaling within a single architecture. Beyond the empirical results, this work reframes multi-model systems as objects governed by regular and predictable scaling behavior rather than as ad-hoc engineering solutions. The proposed law identifies model diversity as a fundamental scaling dimension, alongside parameters, data, and compute. From this perspective, the perceived limits of single-model scaling appear to be conditional on architectural lineage rather than absolute boundaries. While the Law of Multi-model Collaboration defines theoretical upper bounds rather than realizable systems, such bounds are essential for clarifying what is achievable in principle. We hope this work inspires further research into diversity-aware scaling, principled integration mechanisms, and broader theories of collective intelligence in machine learning.